In a particular community, it is claimed that the mean household water usage for a particular month is 45 cubic meters. The following year, a countrywide water conservation campaign was conducted. Forty nine homes were randomly selected and found that the mean consumption is 54 cubic meters with a standard deviation of 5 cubic meters. Is there enough evidence to say that the mean household water usage per month is higher than 45 cubic meters at a=0.01?


1
Expert's answer
2022-03-19T02:41:58-0400


Null Hypothesis, "H_o:\mu>45" , There is enough evidence to support the claim.

Alternative Hypothesis, "H_a: μ \le45" ,There is no enough evidence to support the claim.


Test statistic,


"z= \dfrac{(\bar{x} - \mu)}{(\frac{\sigma}{\sqrt{n})}}

\\[9pt]

z = \dfrac{(45 - 54)}{(\frac{5}{\sqrt{49}}}
\\[9pt]


z = -12.6"


P-value Approach


P-value "=P(z>-12.6)=1"


As "P-value \ge 0.01" , we Accept null hypothesis.


Hence , There is enough evidence to support the claim that the mean household water usage for a particular month is greater than 45 cubic meters.